# Black Hole Cold Brew: Fermi Degeneracy Pressure

### v1

@@ VI Conclusions
Acknowledgements
WXF was supported by Tsinghua’s Shuimu Scholar Fellowship and the China Postdoctoral Science Foundation under Grant No. 2024M761594. HBY was supported by the U.S. Department of Energy under Grant No. DE-SC0008541 and the John Templeton Foundation under Grant 63599. YMZ was supported by the GRF grant 11302824 from the Hong Kong Research Grants Council. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ Appendix C Numerical results of the truncated Fermi–Dirac model
In Tables. 17, we show marginally stable configurations for $\alpha(R)$ ranging from 50-50 to $1$. These configurations satisfy the adiabatic index criterion $\langle\gamma\rangle=\gamma_{{\rm cr}}$. From the 1st to 12th columns, we show their boundary temperature $b=k_{B}T(R)/mc^{2}$, central cutoff function $w(0)=\epsilon_{c}(0)/k_{B}T(0)$, gravitational mass $\hat{M}$, system radius $\hat{R}$, compactness $C=GM(R)/c^{2}R=\hat{M}/\hat{R}$, central interior redshift $Z(0)=e^{-\Phi(0)}-1$, central energy cutoff $\epsilon_{c}(0)$, central energy density $\hat{\rho}(0)$, central pressure $\hat{p}(0)$, central velocity dispersion $v(0)$, boundary velocity dispersion $v(R)$, and $\langle\gamma\rangle=\gamma_{\text{cr}}$.
Table 1: Marginally stable configurations at different boundary temperature $b=k_{B}T(R)/mc^{2}$ for a Fermi–-Dirac distribution with $\alpha(R)=-50$, corresponding to a non-degenerate core. In this case, dynamical instability cannot be achieved for $b<0.09$. Marginally stable configurations are indistinguishable for a Fermi–Dirac distribution with $\alpha(R)=-50$ and a Maxwell–Boltzmann distribution for $b\geq 0.09$.

### v2

@@ VI Conclusions
Acknowledgments
WXF was supported by Tsinghua’s Shuimu Scholar Fellowship and the China Postdoctoral Science Foundation under Grant No. 2024M761594. HBY was supported by the U.S. Department of Energy under Grant No. DE-SC0008541 and the John Templeton Foundation under Grant No. 63599. YZ was supported by the GRF Grants No. 11302824 and No. 11310925 from the Research Grants Council, University Grants Committee, and the Grants No. 9610645 and No. 7020130 from the City University of Hong Kong. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ Appendix C Numerical results of the truncated Fermi–Dirac model
In Tables 17, we show marginally stable configurations for $\alpha(R)$ ranging from 50-50 to $1$. These configurations satisfy the adiabatic index criterion $\langle\gamma\rangle=\gamma_{{\rm cr}}$. From the 1st to 12th columns, we show their boundary temperature $b=k_{B}T(R)/mc^{2}$, central cutoff function $w(0)=\epsilon_{c}(0)/k_{B}T(0)$, gravitational mass $\hat{M}$, system radius $\hat{R}$, compactness $C=GM(R)/c^{2}R=\hat{M}/\hat{R}$, central interior redshift $Z(0)=e^{-\Phi(0)}-1$, central energy cutoff $\epsilon_{c}(0)$, central energy density $\hat{\rho}(0)$, central pressure $\hat{p}(0)$, central velocity dispersion $v(0)$, boundary velocity dispersion $v(R)$, and $\langle\gamma\rangle=\gamma_{\text{cr}}$.
Table 1: Marginally stable configurations at different boundary temperature $b=k_{B}T(R)/mc^{2}$ for a Fermi–Dirac distribution with $\alpha(R)=-50$, corresponding to a non-degenerate core. In this case, dynamical instability cannot be achieved for $b<0.09$. Marginally stable configurations are indistinguishable for a Fermi–Dirac distribution with $\alpha(R)=-50$ and a Maxwell–Boltzmann distribution for $b\geq 0.09$.